4NH3+302-> 2N2+6H2O determine the amount of grams of N2 (fw= 28 g/mol) is produced if 4.03 miles of NH3 react


1
Expert's answer
2021-02-22T05:57:13-0500

Quality of moles N2 = 4.03/4*2=2.015. 2.015*28 g/mol= 56.42 g.